# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part II: posterior mean approximation

### v2

@@ Abstract
In this work, we construct optimal low-rank approximations for the Gaussian posterior distribution in linear Gaussian inverse problems with possibly infinite-dimensional separable Hilbert parameter spaces and finite-dimensional data spaces. We consider different approximation families for the posterior. We first consider approximate posteriors in which the means vary among a class of either structure-preserving or structure-ignoring low-rank transformations of the data, and in which the posterior covariance is kept fixed. We give necessary and sufficient conditions for these approximating posteriors to be equivalent to the exact posterior, for all possible realisations of the data simultaneously. For such approximations, we measure approximation error with the Kullback–Leibler, Rényi and Amari $\alpha$-divergences for $\alpha\in(0,1)$, and with the Hellinger distance, all averaged over the data distribution. With these losses, we find the optimal approximations and formulate an equivalent condition for their uniqueness, extending the work in finite dimensions of Spantini et al. (SIAM J. Sci. Comput. 2015). We then consider joint approximation of the mean and covariance, by also varying the posterior covariance over the low-rank updates considered in Part I of this work. For the reverse Kullback–Leibler divergence, we show that the separate optimal approximations of the mean and of the covariance can be combined to yield an optimal joint approximation of the mean and covariance. In addition, we interpret the joint approximation with the optimal structure-ignoring approximate mean in terms of an optimal projector in parameter space.
Keywords: nonparametric linear Bayesian inverse problems, Gaussian measures, low-rank operator approximation, equivalent measure approximation MSC codes: 28C20, 47A58, 60G15, 62F15, 62G05
@@ 1 Introduction
Linear inverse problems are characterised by a linear map $G$ that encodes the underlying model and the observation process of the problem at hand. That is, $G$ describes the known relationship between the unknown parameter $x^{\dagger}$ to be inferred and the data, which is a noisy observation of $Gx^{\dagger}$. The parameter $x^{\dagger}$ is often a function, such as a diffusivity field in a partial differential equation. The Bayesian method for inferring $x^{\dagger}$ involves considering $x^{\dagger}$ as a random variable $X$ with specified distribution and finding the conditional distribution of $X$ given the data. The prior distribution is the chosen distribution of $X$ and the posterior distribution is the resulting conditional distribution of $X$ given the data. The spread of the posterior distribution can then be interpreted as a quantification of uncertainty.
@@
see [34, Example 6.23]. The posterior covariance $\mathcal{C}_{\textup{pos}}$ is independent of $y$; only the posterior mean $m_{\textup{pos}}$ depends on the realisation of the data.
@@
For such finite-dimensional posterior distributions, various works have studied its approximation, which can prove beneficial in terms of computation and storage. The update from prior to posterior distribution is determined by the choice of prior, by the structure (1) of the inverse problem and by the observed data $y$. Low dimensionality of this update lies at the core of approximation procedures considered in [15, 9, 37, 22, 21, 8, 33]. In [15], low-rank approximation for Gaussian linear inverse problems is considered, while [33] proves optimality for low-rank approximations of posterior mean and covariance. Low-rank approximation for nonlinear and non-Gaussian problems is studied in [9, 8, 37, 22, 21]. The work of [8] describes an algorithm which exploits the low-rank structure of the prior-to-posterior update for certain nonlinear problems based on the ideas developed in finite dimensions, but which can also target infinite-dimensional posteriors. A common feature of these approximations is that they exploit the low-rank structure of the Bayesian prior-to-posterior update, and not just low-rank structure of the prior or forward model. Also other approximation methods exist, such as variational methods, e.g. [26].
The optimality of specific low-rank approximations of the posterior mean in finite-dimensional linear Gaussian inverse problems is studied in [33]. In [33, Section 4], the approximation error is quantified by considering a Bayes risk, which averages over the data. A goal-oriented version is constructed in [32]. The approximation method developed in [22] also targets approximation of the posterior distribution, and hence the posterior mean, but does so for a specific realisation of $y$.
Instead of discretising the problem, optimal low-rank approximations can also be studied directly for the infinite-dimensional posterior. In order to show consistency of the optimal low-rank approximations constructed for discretised versions of the inverse problem, an optimal low-rank approximation problem in the infinite-dimensional setting is required. In this work we aim to analyse and provide such optimal low-rank approximations for the posterior mean directly in the Hilbert space formulation. Furthermore, using the results of Part I of this paper in [6] on optimal low-rank posterior covariance approximations, we also identify low-rank joint approximations of the posterior mean and covariance. In this way, we obtain discretisation-independent and dimension-independent optimal low-rank posterior approximations.
@@ 1.2 Contributions
We formulate two types of low-rank posterior mean approximation: structure-preserving and structure-ignoring approximations. One type preserves the structure of the prior-to-posterior mean update as a function of the data, while the other does not. Keeping the exact posterior covariance fixed, the posterior mean approximations lead to approximate posterior distributions. Not every low-rank posterior mean updates retains equivalence between the corresponding approximate posterior distribution and the exact posterior. In fact, direct generalisation to infinite dimensions of the low-rank updates of [33, Section 4] leads to nonequivalent approximations in general. In Proposition 5.5, we characterise, for both the structure-preserving and structure-ignoring posterior mean approximations, which approximations satisfy this equivalence property. Here, equivalence holds not only for one realisation of the data $y$, but for all realisations in a set of probability 1. This is the first main contribution of the paper.
@@
The second main contribution is to solve the Gaussian measure approximation problems for approximating the posterior mean using the low-rank update classes mentioned in the previous paragraph. We keep the exact posterior covariance fixed and quantify the accuracy of an approximation using the Rényi, Amari, Hellinger, and forward and reverse Kullback-Leibler divergences, averaged over the data distribution. That is, we consider approximations of the mean that are accurate on average, rather than for a specific realisation of $y$. These losses are related to the weighted Bayes risk considered in the finite-dimensional case of [33] and are a natural generalisation to infinite dimensions. The approximation problems rely on a generalisation of the result on reduced-rank matrix approximation by [31] and [16] to infinite dimensions, which can be found in [5]. The solutions and the corresponding minimal losses are identified in Theorems 5.10 and 5.11 and Corollary 5.12. The resulting optimal approximations share the property with $m_{\textup{pos}}$ that they lie in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$ with probability 1, and hence in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ with probability 1, since $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ for Gaussian linear inverse problems, see [34, eq. (6.13a)]. Theorems 5.10 and 5.11 and Corollary 5.12 thus extend the results of [33, Section 4] to an infinite-dimensional setting, and also give necessary and sufficient conditions for uniqueness of the optimal approximations.
@@
The third main contribution is to consider the family of measure approximation problems where both the posterior mean and posterior covariance are jointly approximated. We construct approximations of the posterior which are equivalent to the exact posterior, for all realisations of $Y$ in a set of probability 1. We measure the error in terms of the reverse Kullback–Leibler divergence, averaged over $Y$. The reverse Kullback–Leibler divergence is given by $\int\log(\frac{\operatorname{d}\!{\widetilde{\mu}_{\textup{pos}}}}{\operatorname{d}\!{\mu_{\textup{pos}}}})\operatorname{d}\!{\widetilde{\mu}_{\textup{pos}}}$, where $\widetilde{\mu}_{\textup{pos}}$ and $\mu_{\textup{pos}}=\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$ denote the approximate posterior and exact posterior respectively. This divergence is important in variational approximation methods, see e.g.  [27, Theorem 5]. In Proposition 6.1, we exploit the Pythagorean structure of the expression of the Kullback–Leibler divergence between Gaussians. This allows us to show that the problem of finding an optimal low-rank joint approximation of the mean and covariance can be solved by combining an optimal solution of the low-rank covariance approximation problem in part I of this paper—c.f. [6, Theorem 4.21]—with an optimal solution of the low-rank mean approximation problem given in Theorems 5.10 and 5.11 below. The mean, covariance and joint approximation problems have the same necessary and sufficient condition for uniqueness of solutions. The optimal joint approximation result of Proposition 6.1 and its interpretation via optimal projection given in Proposition 7.1 provide a perspective on low-rank posterior Gaussian measure approximation which combines the insights obtained in the separate mean and covariance approximation procedures.
@@
As shown in Part I and recalled below in Proposition 3.4, the Bayesian prior-to-posterior update occurs only on a finite-dimensional subspace of the parameter space. The optimal joint approximation to the posterior only differs significantly from the prior in a few directions of the parameter space, if the optimal approximation is accurate. This follows from Proposition 7.1, which shows that the optimal approximate posterior that results from the structure-ignoring posterior mean approximation can be obtained as the exact posterior corresponding to a projected version of the Bayesian inverse problem (1), in which $G$ is precomposed by a low-rank projector in parameter space. Thus, if the low-rank approximation is accurate, the prior-to-posterior update on the infinite-dimensional parameter space essentially occurs on a low-dimensional subspace of the parameter space.
@@ 1.4 Notation
[…]
@@
For $T\in\mathcal{B}(\mathcal{H})$, we denote by $T^{\dagger}$ the Moore–Penrose inverse of $T$, also known as the generalised inverse and pseudo-inverse of $T$, c.f. [13, Definition 2.2], [12, Section B.2] or [18, Definition 3.5.7]. It holds that $T^{\dagger}$ is bounded if and only if $\operatorname{ran}{T}$ is closed, c.f. [13, Proposition 2.4]. If $T$ is injective, then $T^{\dagger}=T^{-1}$ on $\operatorname{ran}{T}$.
@@
We indicate the replacement of $a$ with $b$ by ‘$a\leftarrow b$’.
@@ 2 Low-rank posterior mean approximations
We have $H=(\mathcal{C}_{\textup{obs}}^{-1/2}G)^{*}(\mathcal{C}_{\textup{obs}}^{-1/2}G)$ and hence $H$ is self-adjoint and nonnegative. Furthermore, by Lemma A.6 and the invertibility of $\mathcal{C}_{\textup{obs}}^{-1/2}$, $\operatorname{rank}\left({H}\right)=\operatorname{rank}\left({(\mathcal{C}_{\textup{obs}}^{-1/2}G)^{*}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{obs}}^{-1/2}G}\right)=\operatorname{rank}\left({G}\right)$.
With the chosen distributions of $X$ and $Y|X$, we have also specified the distributions of $Y$ and $X|Y=y$, i.e. the data distribution and the posterior distribution. They are both Gaussian: $Y\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})$ and $X|Y=y\sim\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$, where by [34, Example 6.23],
@@
Equation (3c) requires some interpretation. Since $\mu_{\textup{pr}}$ is nondegenerate by 2.1, $\operatorname{supp}\mu_{\textup{pr}}=\mathcal{H}$, c.f. [3, Definition 3.6.2] and $\mathcal{C}_{\textup{pr}}$ is positive, hence injective, c.f. Lemmas A.12 and A.2. Therefore, we can invert $\mathcal{C}_{\textup{pr}}$ on its range $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. Also $\mathcal{C}_{\textup{pr}}^{1/2}$ is injective, and hence $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ is dense in $\mathcal{H}$, see Lemmas A.5 and A.4. For a fixed $y$, the measures $\mu_{\textup{pr}}$ and $\mu_{\textup{pos}}(y)$ are equivalent, see [34, Theorem 6.31]. Thus, by the Feldman–Hajek theorem, which is recalled in Theorem 3.2, also $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ is dense in $\mathcal{H}$. We conclude that also $\mathcal{C}_{\textup{pos}}$ and $\mathcal{C}_{\textup{pos}}^{1/2}$ are injective, and $\mathcal{C}_{\textup{pos}}^{-1}$ is a densely-defined operator with $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. Equation (3c) now states that $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+H}$. Since $H=G^{*}\mathcal{C}_{\textup{obs}}^{-1}G\in\mathcal{B}(\mathcal{H})$, c.f. (2), is defined on all of $\mathcal{H}$, this shows $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}}$. Hence, $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, and this subspace forms the domain of definition of (3c).
@@
[…]
In this way, we ensure that by approximating the posterior mean by $Ay$ for $A\in\mathscr{M}^{(i)}_{r}$, the equivalence with $\mu_{\textup{pos}}(y)$ is maintained for all $y$ in a set of probability 1 with respect to the distribution of $Y$. We stress that $A$ is constructed before a specific realisation $y$ of $Y$ is observed. The structure-preserving class in (4a) takes into account properties of the posterior mean and covariance that are implied by (3a)-(3b). In contrast, the structure-ignoring class in (4b) ignores these properties and only requires that the posterior mean is a linear transformation of the data and that the resulting approximate posterior approximation is equivalent to the exact posterior. We note that the rank-$r$ update $-B$ of $\mathcal{C}_{\textup{pr}}$ in (4a) is not required to be self-adjoint. However, as we shall see in Section 5, the posterior mean approximations of the form (4a) which are optimal, in the sense specified in Section 5, do in fact correspond to self-adjoint updates $-B$.
@@
Because the rank of $A$ and $B$ in (4a) and (4b) are restricted and may be much smaller than $n$, we refer to $Ay$ for $A\in\mathscr{M}^{(i)}_{r}$, $i=1,2$, as a ‘low-rank’ approximation of $m_{\textup{pos}}(y)$. If $\dim{\mathcal{H}}<\infty$, then $\mathscr{M}^{(i)}_{r}$ coincides with the approximation classes considered in [33, Section 4].

### v3

# Optimal low-rank posterior mean and distribution approximation in linear Gaussian inverse problems on Hilbert spaces

@@ Abstract
In this work, we construct optimal low-rank approximations for the Gaussian posterior distribution in linear Gaussian inverse problems with possibly infinite-dimensional separable Hilbert parameter spaces and finite-dimensional data spaces. We consider different approximation families for the posterior. We first consider approximate posteriors in which the means vary among a class of either structure-preserving or structure-ignoring low-rank transformations of the data, and in which the posterior covariance is kept fixed. We give necessary and sufficient conditions for these approximating posteriors to be equivalent to the exact posterior, for all possible realisations of the data simultaneously. For such approximations, we measure approximation error with the Kullback–Leibler, Rényi and Amari $\alpha$-divergences for $\alpha\in(0,1)$, and with the Hellinger distance, all averaged over the data distribution. With these losses, we find the optimal approximations and formulate an equivalent condition for their uniqueness, extending the work in finite dimensions of Spantini et al. (SIAM J. Sci. Comput. 2015). We then consider joint low-rank approximation of the mean and covariance. For the reverse Kullback–Leibler divergence, we show that the separate optimal approximations of the mean and of the covariance can be combined to yield an optimal joint approximation of the mean and covariance. In addition, we interpret the joint approximation with the optimal structure-ignoring approximate mean in terms of an optimal projector in parameter space, showing this approximation amounts to solving a Bayesian inverse problem with projected forward model.
Keywords: nonparametric linear Bayesian inverse problems, Gaussian measure, low-rank operator approximation, equivalent measure approximation, projected inverse problem MSC codes: 28C20, 47A58, 60G15, 62F15, 62G05
@@ 1 Introduction
Linear inverse problems are characterised by a linear map $G$ that encodes the underlying model and the observation process of the problem at hand. That is, $G$ describes the known relationship between the unknown parameter $x^{\dagger}$ to be inferred and the data, which is a noisy observation of $Gx^{\dagger}$. The parameter $x^{\dagger}$ is often a function, such as a diffusivity field in a partial differential equation.
Inference on $x^{\dagger}$ essentially amounts to inverting the operator $G$. Such inversion is typically an ill-posed operation, due to the smoothing nature of $G$. For example, if $G$ involves application of an elliptic partial differential equation, then $G$ typically has quickly decaying spectrum, since the inverse Laplacian has quickly decaying spectrum. Furthermore, inference of a function $x^{\dagger}$ based on a finite amount of observations need not be uniquely possible. For these reasons, regularisation is required. Bayesian methods can be seen as a way to regularise the inverse problem, and also naturally allow for uncertainty quantification. To quantify the uncertainty, the posterior covariance operator is essential.
The Bayesian method for inferring $x^{\dagger}$ involves considering $x^{\dagger}$ as a random variable $X$ with specified distribution and finding the conditional distribution of $X$ given the data. The prior distribution is the chosen distribution of $X$ and the posterior distribution is the resulting conditional distribution of $X$ given the data. The spread of the posterior distribution can then be interpreted as a quantification of uncertainty.
@@
see [40, Example 6.23]. The posterior covariance $\mathcal{C}_{\textup{pos}}$ is independent of $y$; only the posterior mean $m_{\textup{pos}}$ depends on the realisation of the data.
@@
For such finite-dimensional posterior distributions, various works have studied its approximation, which for tractability in terms of computation and storage may be essential. The update from prior to posterior distribution is determined by the choice of prior, by the structure (1) of the inverse problem and by the observed data $y$. Low dimensionality of this update lies at the core of approximation procedures considered in [20, 14, 43, 27, 26, 13, 39]. In [20], low-rank approximation for Gaussian linear inverse problems is considered, while [39] proves optimality for low-rank approximations of posterior mean and covariance. Low-rank approximation for nonlinear and non-Gaussian problems is studied in [14, 13, 43, 27, 26]. The work of [13] describes an algorithm which exploits the low-rank structure of the prior-to-posterior update for certain nonlinear problems based on the ideas developed in finite dimensions, but which can also target infinite-dimensional posteriors. A common feature of these approximations is that they exploit the low-rank structure of the Bayesian prior-to-posterior update, and not just low-rank structure of the prior or forward model. Also other approximation methods exist, such as variational methods, e.g. [31].
The optimality of specific low-rank approximations of the posterior mean in finite-dimensional linear Gaussian inverse problems is studied in [39]. Such an approximation may prove useful in a many-query setting, in which the posterior mean has to be recomputed for many different realisations of the data. In [39, Section 4], the approximation error is quantified by considering a Bayes risk, which averages over the data. A goal-oriented version is constructed in [38]. The approximation method developed in [27] also targets approximation of the posterior distribution, and hence the posterior mean, but does so for a specific realisation of $y$.
Instead of discretising the problem, optimal low-rank approximations can also be studied directly for the infinite-dimensional posterior. In order to show consistency of the optimal low-rank approximations constructed for discretised versions of the inverse problem, an optimal low-rank approximation problem in the infinite-dimensional setting is required. Then, once a specific approximation scheme is chosen for a given inverse problem, this infinite-dimensional optimal approximation can be used to show discretisation independence of the approximation method. This is similar in spirit to how [8] shows dimension independence of a sampling scheme for a finite element based discretision of certain partial differential equations, using the infinite-dimensional results on sampling methods established in [12, 3]. Numerical evidence that discretisation independence should hold in a specific setting was found in [7].
In this work we aim to analyse and provide such optimal low-rank approximations for the posterior mean directly in the Hilbert space formulation. Furthermore, using the results of [10] on optimal low-rank posterior covariance approximations, we also identify low-rank joint approximations of the posterior mean and covariance. This allows us to obtain discretisation-independent and dimension-independent optimal low-rank posterior approximations.
@@ 1.2 Contributions
We formulate two types of low-rank posterior mean approximation: structure-preserving and structure-ignoring approximations. One type preserves the structure of the prior-to-posterior mean update as a function of the data, while the other does not. Keeping the exact posterior covariance fixed, the posterior mean approximations lead to approximate posterior distributions. Not every low-rank posterior mean updates retains equivalence between the corresponding approximate posterior distribution and the exact posterior. In fact, direct generalisation to infinite dimensions of the low-rank updates of [39, Section 4] leads to nonequivalent approximations in general. In Proposition 5.5, we characterise, for both the structure-preserving and structure-ignoring posterior mean approximations, which approximations satisfy this equivalence property. Here, equivalence holds not only for one realisation of the data $y$, but for all realisations in a set of probability 1. This is the first main contribution of the paper.
@@
The second main contribution is to solve the Gaussian measure approximation problems for approximating the posterior mean using the low-rank update classes mentioned in the previous paragraph. We keep the exact posterior covariance fixed and quantify the accuracy of an approximation using the Rényi, Amari, Hellinger, and forward and reverse Kullback-Leibler divergences, averaged over the data distribution. That is, we consider approximations of the mean that are accurate on average, rather than for a specific realisation of $y$. These losses are related to the weighted Bayes risk considered in the finite-dimensional case of [39] and are a natural generalisation to infinite dimensions. The approximation problems rely on a generalisation of the result on reduced-rank matrix approximation by [37] and [21] to infinite dimensions, which can be found in [9]. The solutions and the corresponding minimal losses are identified in Theorems 5.10 and 5.11 and Corollary 5.12. The resulting optimal approximations share the property with $m_{\textup{pos}}$ that they lie in $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$ with probability 1, and hence in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ with probability 1, since $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ for Gaussian linear inverse problems, see [40, eq. (6.13a)]. Theorems 5.10 and 5.11 and Corollary 5.12 thus extend the results of [39, Section 4] to an infinite-dimensional setting, and also give necessary and sufficient conditions for uniqueness of the optimal approximations.
@@
The third main contribution is to consider the family of measure approximation problems where both the posterior mean and posterior covariance are jointly approximated. We construct approximations of the posterior which are equivalent to the exact posterior, for all realisations of $Y$ in a set of probability 1. We measure the error in terms of the reverse Kullback–Leibler divergence, averaged over $Y$. The reverse Kullback–Leibler divergence is given by $\int\log(\frac{\operatorname{d}\!{\widetilde{\mu}_{\textup{pos}}}}{\operatorname{d}\!{\mu_{\textup{pos}}}})\operatorname{d}\!{\widetilde{\mu}_{\textup{pos}}}$, where $\widetilde{\mu}_{\textup{pos}}$ and $\mu_{\textup{pos}}=\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$ denote the approximate posterior and exact posterior respectively. This divergence is important in variational approximation methods, see e.g.  [32, Theorem 5]. In Proposition 6.1, we exploit the Pythagorean structure of the expression of the Kullback–Leibler divergence between Gaussians. This allows us to show that the problem of finding an optimal low-rank joint approximation of the mean and covariance can be solved by combining an optimal solution of the low-rank covariance approximation problem in [10, Theorem 4.21] with an optimal solution of the low-rank mean approximation problem given in Theorems 5.10 and 5.11 below. The mean, covariance and joint approximation problems have the same necessary and sufficient condition for uniqueness of solutions. The optimal joint approximation result of Proposition 6.1 and its interpretation via optimal projection given in Proposition 7.1 provide a perspective on low-rank posterior Gaussian measure approximation which combines the insights obtained in the separate mean and covariance approximation procedures.
@@
As shown in [10] and recalled below in Proposition 3.4, the Bayesian prior-to-posterior update occurs only on a finite-dimensional subspace of the parameter space. The optimal joint approximation to the posterior only differs significantly from the prior in a few directions of the parameter space, if the optimal approximation is accurate. This follows from Proposition 7.1, which shows that the optimal approximate posterior that results from the structure-ignoring posterior mean approximation can be obtained as the exact posterior corresponding to a projected version of the Bayesian inverse problem (1), in which $G$ is precomposed by a low-rank projector in parameter space. Thus, if the low-rank approximation is accurate, the prior-to-posterior update on the infinite-dimensional parameter space essentially occurs on a low-dimensional subspace of the parameter space.
@@ 1.4 Notation
To introduce the notation, we let $\mathcal{H}$ and $\mathcal{K}$ be separable Hilbert spaces, that is, complete inner product spaces with a countable orthonormal basis (ONB). We denote the linear spaces of linear operators defined with domain $\mathcal{H}$ and codomain $\mathcal{K}$ which are bounded, compact and finite-rank by, respectively, $\mathcal{B}(\mathcal{H},\mathcal{K})$, $\mathcal{B}_{0}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}_{00}(\mathcal{H},\mathcal{K})$. A linear operator is said to have ‘finite rank’ if it is bounded and its range is finite-dimensional. The set of finite-rank operators which have rank at most $r\in\mathbb{N}$ is denoted by $\mathcal{B}_{00,r}(\mathcal{H},\mathcal{K})$. The above sets are all endowed with the operator norm $\lVert\cdot\rVert$ defined by $\lVert T\rVert\coloneqq\sup\{\lVert Th\rVert:\lVert h\rVert\leq 1\}$. The trace-class and Hilbert–Schmidt operators are compact operators with summable and square-summable eigenvalue sequence respectively, and are denoted by $L_{1}(\mathcal{H},\mathcal{K})$ and $L_{2}(\mathcal{H},\mathcal{K})$ respectively. Their respective norms are denoted by $\lVert\cdot\rVert_{L_{1}(\mathcal{H},\mathcal{K})}$ and $\lVert\cdot\rVert_{L_{2}(\mathcal{H},\mathcal{K})}$. Thus, $\lVert T\rVert_{L_{1}(\mathcal{H},\mathcal{K})}$ and $\lVert T\rVert_{L_{2}(\mathcal{H},\mathcal{K})}^{2}$ are computed by summing respectively the absolute values and squares of the singular values of $T$. If $\mathcal{H}=\mathcal{K}$, then we write $\mathcal{B}(\mathcal{H})$ instead of $\mathcal{B}(\mathcal{H},\mathcal{K})$, and similarly for the other sets above. We have the inclusion of sets $\mathcal{B}_{00,r}(\mathcal{H})\subset\mathcal{B}_{00}(\mathcal{H})\subset L_{1}(\mathcal{H})\subset L_{2}(\mathcal{H})\subset\mathcal{B}_{0}(\mathcal{H})\subset\mathcal{B}(\mathcal{H})$.
@@
For $T\in\mathcal{B}(\mathcal{H})$, we denote by $T^{\dagger}$ the Moore–Penrose inverse of $T$, also known as the generalised inverse and pseudo-inverse of $T$, c.f. [18, Definition 2.2], [17, Section B.2] or [23, Definition 3.5.7]. It holds that $T^{\dagger}$ is bounded if and only if $\operatorname{ran}{T}$ is closed, c.f. [18, Proposition 2.4]. If $T$ is injective, then $T^{\dagger}=T^{-1}$ on $\operatorname{ran}{T}$.
@@
We indicate the replacement of $a$ with $b$ by $a\leftarrow b$’.
@@ 2 Low-rank posterior mean approximations
In statistics, $H$ is also known as the Fisher information operator, but we shall refer to it as “the Hessian”. We have $H=(\mathcal{C}_{\textup{obs}}^{-1/2}G)^{*}(\mathcal{C}_{\textup{obs}}^{-1/2}G)$ and hence $H$ is self-adjoint and nonnegative. Furthermore, by Lemma A.6 and the invertibility of $\mathcal{C}_{\textup{obs}}^{-1/2}$, $\operatorname{rank}\left({H}\right)=\operatorname{rank}\left({(\mathcal{C}_{\textup{obs}}^{-1/2}G)^{*}}\right)=\operatorname{rank}\left({\mathcal{C}_{\textup{obs}}^{-1/2}G}\right)=\operatorname{rank}\left({G}\right)$.
With the chosen distributions of $X$ and $Y|X$, we have also specified the distributions of $Y$ and $X|Y=y$, i.e. the data distribution and the posterior distribution. They are both Gaussian: $Y\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})$ and $X|Y=y\sim\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pos}})$, where by [40, Example 6.23],
@@
Equation (3c) requires some interpretation. Since $\mu_{\textup{pr}}$ is nondegenerate by 2.1, $\operatorname{supp}\mu_{\textup{pr}}=\mathcal{H}$, c.f. [4, Definition 3.6.2] and $\mathcal{C}_{\textup{pr}}$ is positive, hence injective, c.f. Lemmas A.12 and A.2. Therefore, we can invert $\mathcal{C}_{\textup{pr}}$ on its range $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. Also $\mathcal{C}_{\textup{pr}}^{1/2}$ is injective, and hence $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ is dense in $\mathcal{H}$, see Lemmas A.5 and A.4. For a fixed $y$, the measures $\mu_{\textup{pr}}$ and $\mu_{\textup{pos}}(y)$ are equivalent, see [40, Theorem 6.31]. Thus, by the Feldman–Hajek theorem, which is recalled in Theorem 3.2, also $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ is dense in $\mathcal{H}$. We conclude that also $\mathcal{C}_{\textup{pos}}$ and $\mathcal{C}_{\textup{pos}}^{1/2}$ are injective, and $\mathcal{C}_{\textup{pos}}^{-1}$ is a densely-defined operator with $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. Equation (3c) now states that $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+H}$. Since $H=G^{*}\mathcal{C}_{\textup{obs}}^{-1}G\in\mathcal{B}(\mathcal{H})$, c.f. (2), is defined on all of $\mathcal{H}$, this shows $\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}}$. Hence, $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, and this subspace forms the domain of definition of (3c).
@@
$\mathscr{M}^{(1)}_{r}\coloneqq$ $\{(\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ B\in\mathcal{B}_{00,r}(\mathcal{H}),\ \mathcal{N}((\mathcal{C}_{\textup{pr}}-B)G^{*}\mathcal{C}_{\textup{obs}}^{-1}Y,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}(Y)\quad\text{a.s.}\},$ (4a)
$\mathscr{M}^{(2)}_{r}\coloneqq$ $\{A\in\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}):\ \mathcal{N}(AY,\mathcal{C}_{\textup{pos}})\sim\mu_{\textup{pos}}(Y)\quad\text{a.s.}\}.$ (4b)
In this way, we ensure that by approximating the posterior mean by $Ay$ for $A\in\mathscr{M}^{(i)}_{r}$, the equivalence with $\mu_{\textup{pos}}(y)$ is maintained for all $y$ in a set of probability 1 with respect to the distribution of $Y$. We stress that $A$ is constructed before a specific realisation $y$ of $Y$ is observed. The structure-preserving class in (4a) takes into account properties of the posterior mean and covariance that are implied by (3a)-(3b). In contrast, the structure-ignoring class in (4b) ignores these properties and only requires that the posterior mean is a linear transformation of the data and that the resulting approximate posterior approximation is equivalent to the exact posterior. We note that the rank-$r$ update B-B of $\mathcal{C}_{\textup{pr}}$ in (4a) is not required to be self-adjoint. However, as we shall see in Section 5, the posterior mean approximations of the form (4a) which are optimal, in the sense specified in Section 5, do in fact correspond to self-adjoint updates B-B.
@@
Because the rank of $A$ and $B$ in (4a) and (4b) are restricted and may be much smaller than $n$, we refer to AyAy for $A\in\mathscr{M}^{(i)}_{r}$, $i=1,2$, as a ‘low-rank’ approximation of $m_{\textup{pos}}(y)$. If $\dim{\mathcal{H}}<\infty$, then $\mathscr{M}^{(i)}_{r}$ coincides with the approximation classes considered in [39, Section 4].